if the 4 th term of a h is 1/3 and 7 th term is 1/n then 6 th term is what ?

if the 4th term of a h is 1/3 and 7th term is 1/n then 6th term is what ?

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
By what i understand by your representation.
I’m assuming you have an HP.
\\a_1,a_2,a_3,a_4,a_5....... \\1/a_i$ is an AP$ \\=>1/a_1+3d=1/3 \\=>1/a_1+6d=1/n \\=>3d=1/n-1/3 =>1/a_7=1/a_1+5d=1/n-{1/n - 1/3 \over 3}
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty